OpenStudy (anonymous):

Help please!! Find the exact value of the radical expression in simplest form. the square root of p to the sixth power + the square root of 9 times p to the sixth power + 3 the square root of p to the sixth power − the square root of p squared 13 times p cubed times the square root of p − 2 times p times the square root of p 13 times p cubed times the square root of p − 2p 7p^3 7p^3 − p

OpenStudy (anonymous):

Can you draw or input the equation?

OpenStudy (anonymous):

\[\sqrt{p^6} + \sqrt{9p^6} + 3 \sqrt{p^6} - \sqrt{p^2}\]

OpenStudy (anonymous):

can you help me? @FriedRice14

OpenStudy (anonymous):

guess not. k

OpenStudy (cwrw238):

sqrt p^6 = p^3 sqrt (9p^6) = 3p^3 so the first 3 terms = p^3 + 3p^3 + 3p^3 = 7p^3 and sqrt p^2 = p

OpenStudy (anonymous):

so it's c

OpenStudy (cwrw238):

no d 7p^3 - p

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

can you please help me with more

OpenStudy (cwrw238):

sorry gotta go